# PoisonedRAG: Knowledge Poisoning Attacks to Retrieval-Augmented Generation of Large Language Models

### v1

@@ 5.3.1 Impact of Hyperparameters in RAG
Figure 5: Impact of the length of $I$ for PoisonedRAG on NQ. Figures 1617 (in Appendix) show results of other datasets.
@@
Impact of similarity metric: Table 7 shows the results when we use different similarity metrics to calculate the similarity of embedding vectors when retrieving texts from a knowledge database for a question. We find that PoisonedRAG achieves similar results for different similarity metrics in both settings.
Impact of LLMs: Table 1 also shows the results of PoisonedRAG for different LLMs in RAG. We find that PoisonedRAG consistently achieves high ASRs.
@@ 5.3.2 Impact of Hyperparameters in PoisonedRAG
Impact of $N$: Figure 4 shows the impact of $N$. We have the following observations. First, ASR increases as $N$ increases when $N\leq k$ ($k=5$ by default). The reason is that more poisoned texts are injected for each target question when $N$ is larger, and thus the retrieved texts for the target question contain more poisoned ones, e.g., Precision increases as $N$ increases and Recall is consistently high. When $N>k$, ASR (or Precision) becomes stable and is consistently high. We note that Recall decreases as $N$ increases when $N>k$. The reason is that at most $k$ poisoned texts could be retrieved. F1-Score measures a tradeoff between Precision and Recall, which first increases and then decreases.
Impact of length $V$ in generating $I$:  To achieve the effectiveness condition, we use a LLM to generate $I$ with length $V$ (a hyperparameter) such that RAG would generate an attacker-chosen target answer for a target question. We study the impact of $V$ on the effectiveness of PoisonedRAG. Figure 5 shows the experimental results. We find that PoisonedRAG achieves similar ASR, Precision, Recall, and F1-Score, which means PoisonedRAG is insensitive to $V$.
Impact of the number of trials $L$ in generating $I$: Figure 6 shows the impact of number of trials $L$ on PoisonedRAG for NQ. We find that PoisonedRAG could achieve high ASRs even when $L=1$ (i.e., one trial is made). As $L$ increases, the ASR first increases and then becomes saturated when $L\geq 10$. Our experimental results demonstrate that a small $L$ (i.e., 10) is sufficient for PoisonedRAG to achieve high ASRs.
Impact of concatenation order of $S$ and $I$:  By default, we concatenate $S$ and $I$ as $S\oplus I$ to craft a poisoned text. We study whether the concatenation order of $S$ and $I$ would influence the effectiveness of PoisonedRAG. Table 8 shows the experimental results, which demonstrate that PoisonedRAG is also effective when we change their order.
@@
Impact of adversarial text generation methods:  In the white-box setting, PoisonedRAG can utilize any existing adversarial text generation methods [75, 76] to optimize $S$ in Equation 6. By default, we use HotFlip [75]. Here we also evaluate the effectiveness of PoisonedRAG when using TextFooler [76], which replaces words with their synonyms to keep semantics meaning. Table 9 shows the results. We find that PoisonedRAG could achieve very high ASR and F1-Score, demonstrating that PoisonedRAG is effective with different adversarial example methods.
@@ 6 Defenses
Many defenses [87, 88, 89, 90, 91, 92, 93, 94, 95, 96] were proposed to defend against data poisoning attacks that compromise the training dataset of a machine learning model. However, most of those defenses are not applicable to defend against PoisonedRAG since PoisonedRAG does not compromise the training dataset of a LLM. Thus, we generalize some widely used defenses against attacks [42, 43, 44] to LLM to defend against PoisonedRAG.
@@ 6.1 Paraphrasing
Paraphrasing [42] was used to defend against prompt injection attacks [40, 46, 48, 49] and jailbreaking attacks [50, 51, 52, 53, 54, 55] to LLMs. We extend paraphrasing to defend against PoisonedRAG. In particular, given a text, the paraphrasing defense utilizes a LLM to paraphrase it. In our scenario, given a question, we use a LLM to paraphrase it before retrieving relevant texts from the knowledge database to generate an answer for it. Recall that PoisonedRAG crafts poisoned texts such that they could be retrieved for a target question. For instance, in the black-box setting, PoisonedRAG prepends the target question to a text $I$ to craft a poisoned text. In the white-box setting, PoisonedRAG optimizes a poisoned text such that a retriever produces similar feature vectors for the poisoned text and the target question. Our insight is that paraphrasing the target question would change its structure. For instance, when the target question is “Who is the CEO of OpenAI?”. The paraphrased question could be “Who holds the position of Chief Executive Officer at OpenAI?”. As a result, poisoned texts may not be retrieved for the paraphrased target question. Note that we do not paraphrase texts in the knowledge database due to high computational costs.
We conduct experiments to evaluate the effectiveness of paraphrasing defense. In particular, for each target question, we generate 5 paraphrased target questions using GPT-4, where the prompt can be found in Appendix E. For each paraphrased target question, we retrieve $k$ texts from the poisoned knowledge database (the poisoned texts are crafted for the original target questions using PoisonedRAG). Then, we generate an answer for the paraphrased target question based on the $k$ retrieved texts. We adopt the same default setting as that in Section 5 (e.g., $k=5$ and 5 injected poisoned texts for each target question). We report the ASR and F1-Score (note that Precision and Recall are the same as F1-Score under our default setting). ASR measures the fraction of paraphrased target questions whose answers are the corresponding attacker-chosen target answers. F1-Score is higher when more poisoned texts designed for a target question are retrieved for the corresponding paraphrased target questions. Table 10 shows our experimental results. We find that PoisonedRAG could still achieve high ASRs and F1-Score, which means paraphrasing defense cannot effectively defend against PoisonedRAG.
@@
Figure 7: The ROC curves for PPL detection defense. The dataset is NQ. The results for other two datasets are in Figures 8 and 9 in Appendix.
@@ 6.2 Perplexity-based Detection
Perplexity (PPL) [97] is widely used to measure the quality of texts, which is also utilized to defend against attacks to LLMs [42, 43, 44]. In particular, given a text, a large perplexity means the given text has a low quality. We utilize perplexity to detect poisoned texts. For instance, in the white-box setting, PoisonedRAG utilizes adversarial attacks to craft poisoned samples, which may influence the quality of poisoned texts. Thus, a text with lower text quality (i.e., high perplexity) is more likely to be poisoned. We calculate the perplexity for all clean texts in the database as well as all poisoned texts crafted by PoisonedRAG. In our experiment, we use the cl100k_base model from OpenAI tiktoken [98] to calculate perplexity.
Figure 7 shows the ROC curve as well as AUC. We find that the false positive rate (FPR) is also very large when the true positive rate (TPR) is very large. This means a large fraction of clean texts are also detected as poisoned texts when poisoned texts are detected. In other words, the perplexity values of poisoned texts are not statistically higher than those of clean texts, which means it is very challenging to detect poisoned texts using perplexity. We suspect the reasons are as follows. Recall that each poisoned text $P$ is the concatenation of $S$ and $I$, i.e., $P=S\oplus I$. The sub-text $I$ is generated by GPT-4, which is of high quality. For PoisonedRAG in the black-box setting, $S$ is the target question, which is a normal text. As a result, the text quality of the poisoned text is normal. We find that the AUC of PoisonedRAG in the white-box setting is slightly larger than that in the black-box setting, which means the text quality is influenced by the optimization but not substantially.
@@
6.3 Duplicate Text Filtering
PoisonedRAG generates each poisoned text independently in both black-box and white-box settings. As a result, it is possible that some poisoned texts could be the same. Thus, we could filter those duplicate texts to defend against PoisonedRAG. We add experiments to filter duplicate texts under the default setting in Section 5. In particular, we calculate the hash value (using the SHA-256 hash function) for each text in a poisoned knowledge database and remove texts with the same hash value. Table 11 compares the ASR with and without defense. We find that the ASR is the same, which means duplicate text filtering cannot successfully filter poisoned texts. The reason is that the sub-text $I$ (generated by GPT-4 in our experiment) in each poisoned text is different, resulting in diverse poisoned texts.
@@ 6.4 Knowledge Expansion
PoisonedRAG injects at most $N$ poisoned texts into a knowledge database for each target question. Thus, if we retrieve $k$ texts, with $k>N$, then it is very likely that $k-N$ texts would be clean ones. This inspires us to retrieve more texts to defend against PoisonedRAG. We call this defense Knowledge Expansion. We conduct experiments under the default setting, where $N=5$. Figures 21, 22, 23 (in Appendix) shows the ASRs, Precision, Recall, and F1-Score for large $k$. We find that this defense still cannot completely defend against our PoisonedRAG even if $k=50$ (around 10% retrieved texts are poisoned ones when injecting $N=5$ poisoned texts for each target question). For instance, PoisonedRAG could still achieve 41% (black-box) and 43% (white-box) ASR on HotpotQA when $k=50$. Additionally, we find that ASR further increases as $N$ increases (shown in Figures 24, 25, 26 in Appendix), which means this defense is less effective when an attacker could inject more poisoned texts into the knowledge database. We note that this defense also incurs large computation costs for a LLM to generate an answer due to the long context (caused by more retrieved texts).

### v2

@@ 5.3.1 Impact of Hyperparameters in RAG
Impact of similarity metric. Table 6 shows the results when we use different similarity metrics to calculate the similarity of embedding vectors when retrieving texts from a knowledge database for a question. We find that PoisonedRAG achieves similar results for different similarity metrics in both settings.
Impact of LLMs. Table 1 also shows the results of PoisonedRAG for different LLMs in RAG. We find that PoisonedRAG consistently achieves high ASRs. We also study the impact of the temperature hyperparameter of the LLM in RAG on PoisonedRAG. Table 18 in Appendix shows the results when setting a large temperature, which demonstrate that the effectiveness of PoisonedRAG is unaffected by the randomness in the decoding process of the LLM.
@@ 5.3.2 Impact of Hyperparameters in PoisonedRAG
Impact of $N$. Figure 4 shows the impact of $N$. We have the following observations. First, ASR increases as $N$ increases when $N\leq k$ ($k=5$ by default). The reason is that more malicious texts are injected for each target question when $N$ is larger, and thus the retrieved texts for the target question contain more malicious ones, e.g., Precision increases as $N$ increases and Recall is consistently high. When $N>k$, ASR (or Precision) becomes stable and is consistently high. We note that Recall decreases as $N$ increases when $N>k$. The reason is that at most $k$ malicious texts could be retrieved. F1-Score measures a tradeoff between Precision and Recall, which first increases and then decreases.
Impact of length $V$ in generating $I$.  To achieve the generation condition, we use an LLM to generate $I$ with length $V$ (a hyperparameter) such that RAG would generate an attacker-chosen target answer for a target question. We study the impact of $V$ on the effectiveness of PoisonedRAG. Figure 15 - 17 (in Appendix) shows the experimental results. We find that PoisonedRAG achieves similar ASR, Precision, Recall, and F1-Score, which means PoisonedRAG is insensitive to $V$.
Impact of the number of trials $L$ in generating $I$. Figure 5 shows the impact of number of trials $L$ on PoisonedRAG for NQ. We find that PoisonedRAG could achieve high ASRs even when $L=1$ (i.e., one trial is made). As $L$ increases, the ASR first increases and then becomes saturated when $L\geq 10$. Our experimental results demonstrate that a small $L$ (i.e., 10) is sufficient for PoisonedRAG to achieve high ASRs.
Impact of concatenation order of $S$ and $I$.  By default, we concatenate $S$ and $I$ as $S\oplus I$ to craft a malicious text. We study whether the concatenation order of $S$ and $I$ would influence the effectiveness of PoisonedRAG. Table 7 shows the experimental results, which demonstrate that PoisonedRAG is also effective when we change their order.
The effectiveness of each attack component.  The effectiveness of our PoisonedRAG depends on 1) whether malicious texts are retrieved, and 2) whether the retrieved malicious texts can make the LLM in RAG generate target answers. We independently evaluate the effectiveness of each part. Table 16 (in Appendix) shows the first part results, demonstrating most malicious texts are retrieved for corresponding target questions. To study the second part, we vary the number of malicious texts (randomly selected) in the $k$ retrieved texts. Table  17 (in Appendix) shows the results. As we can see, the ASR increases as as the number of malicious texts increases. When the number of malicious texts is small, the ASR does not reach 100%. We suspect the reason is that most of the $k$ ($k=5$ by default) retrieved texts are clean ones, making the attack less effective, i.e., the LLM could still generate answers based on clean texts for some target questions. By contrast, when the number of malicious texts is larger than 3 (most of the $k$ retrieved texts are malicious ones), the ASR is very high.
@@
Table 9: The effectiveness of PoisonedRAG when using less powerful LLMs to generate $I$ to achieve generation condition.
Dataset Attack PaLM 2 GPT-3.5 LLaMa-2-7B Vicuna-7B
ASR F1-Score ASR F1-Score ASR F1-Score ASR F1-Score
NQ PoisonedRAG (Black-Box) 0.99 0.97 0.98 0.95 0.92 0.93 0.95 0.95
PoisonedRAG (White-Box) 0.97 1.00 0.98 0.99 0.91 0.98 0.97 0.99
HotpotQA PoisonedRAG (Black-Box) 0.97 1.00 0.99 1.00 0.96 0.99 0.98 1.00
PoisonedRAG (White-Box) 0.96 1.00 0.96 1.00 0.97 1.00 0.99 1.00
MS-MARCO PoisonedRAG (Black-Box) 0.91 0.88 0.89 0.84 0.79 0.80 0.90 0.82
PoisonedRAG (White-Box) 0.95 0.97 0.89 0.95 0.84 0.89 0.92 0.95
Table 10: The effectiveness of PoisonedRAG under advanced RAG.
Dataset Attack Self-RAG [31] CRAG [93]
ASR F1-Score ASR F1-Score
NQ PoisonedRAG (Black-Box) 0.77 0.96 0.78 0.96
PoisonedRAG (White-Box) 0.74 1.0 0.82 1.0
Hotpot QA PoisonedRAG (Black-Box) 0.87 1.0 0.76 1.0
PoisonedRAG (White-Box) 0.79 1.0 0.70 1.0
MS- MARCO PoisonedRAG (Black-Box) 0.73 0.89 0.74 0.89
PoisonedRAG (White-Box) 0.75 0.94 0.72 0.94
@@
Impact of adversarial text generation methods in generating $S$ to achieve retrieval condition.  In the white-box setting, PoisonedRAG can utilize any existing adversarial text generation methods [78, 80] to optimize $S$ in Equation 5. By default, we use HotFlip [78]. Here we also evaluate the effectiveness of PoisonedRAG when using TextFooler [80], which replaces words with their synonyms to keep semantic meanings. Table 8 shows the results, which demonstrate that PoisonedRAG could achieve very high ASR and F1-Score using both methods. We also compare the computational overhead of the two methods in Table 22 (in Appendix). We find that PoisonedRAG could incur higher computational overhead using TextFooler. The reason is that TextFooler aims to keep the semantic meaning when crafting adversarial texts (e.g., using synonyms of words for replacement in optimizing adversarial text). As a result, the candidate word space in each iteration is smaller, which means more iterations are needed for optimization, resulting in higher overhead. However, the adversarial text crafted by TextFooler is more stealthy as it keeps semantics. Our results demonstrate that there is a tradeoff between computational overhead and stealthiness.
Impact of the LLM in generating $I$ to achieve generation condition.  By default, we use GPT-4 to generate $I$ to achieve the generation condition because it is very powerful. We also evaluate whether PoisonedRAG could be effective when using less powerful LLMs to generate $I$. As those LLMs are less powerful, we utilize in-context learning [1] to improve the performance (we provide two demonstration samples to the LLM, please see Appendix L for details). Table 9 shows the experimental results under the default setting, which show our PoisonedRAG is also effective when using less powerful LLMs to generate $I$ with in-context learning.
6 Evaluation for Real-world Applications
We evaluate PoisonedRAG for more sophisticated RAG schemes and two real-world applications, including Wikipedia-based ChatBot and LLM agents.
6.1 Advanced RAG Schemes
In our above experiment, we mainly focus on basic RAG. However, the basic RAG scheme might be insufficient for more complex real-world applications. To this end, many advanced RAG schemes [31, 93, 94, 95, 96] were proposed to improve the performance of the basic RAG scheme. For example, Asai et al. [31] introduced Self-RAG, which trains an LLM that can adaptively use the retrieved contexts on-demand and reflect on its own generations to enhance the factuality and quality of generated answers. Yan et al. [93] proposed CRAG, which uses a lightweight retrieval evaluator to assess the quality (e.g., relevance of retrieved texts to questions) of retrieved contexts, thus enhancing the robustness and correctness of RAG. Roughly speaking, their key idea is to enhance the relevance of the retrieved texts and thus make the LLM more likely to generate correct answers based on the retrieved texts. We conduct experiments to evaluate the effectiveness of PoisonedRAG for these advanced RAG schemes. Table 10 shows PoisonedRAG can achieve high ASRs, demonstrating that those advanced RAG schemes are also vulnerable to our PoisonedRAG. The reason is that the crafted malicious texts are relevant to the target questions, making the LLM generate incorrect answers based on malicious texts.
6.2 Wikipedia-based ChatBot
In our threat model, we consider an attacker can inject malicious texts into a knowledge database collected from Wikipedia by maliciously editing Wikipedia articles [37]. We use a case study to evaluate the effectiveness of PoisonedRAG in this scenario. We used the English Wikipedia dump from Dec. 20, 2018 to create a knowledge database [13]. In particular, each English Wikipedia article (non-text parts are removed) is split into disjoint texts of 100 words. The total number of texts in the knowledge database is 21,015,324. We create a ChatBot based on this knowledge database using the same system prompt as in Appendix B. We evaluate whether our PoisonedRAG is effective for this large knowledge database. We use the default setting of our previous evaluation (in Section 5.1). We reuse the target questions from our previous three datasets (i.e., NQ, HotpotQA, and MS-MARCO) and inject five malicious texts for each target question. Results in Table 11 show PoisonedRAG is effective in this real-world scenario.
Table 11: PoisonedRAG is still effective in a real-world scenario, where the knowledge database consists of 21,015,324 texts from Dec. 20, 2018 Wikipedia dump.
Dataset of Target Questions Attack ASR F1-Score
NQ PoisonedRAG (Black-Box) 0.95 0.95
PoisonedRAG (White-Box) 0.97 0.99
HotpotQA PoisonedRAG (Black-Box) 1.0 1.0
PoisonedRAG (White-Box) 0.94 1.0
MS-MARCO PoisonedRAG (Black-Box) 0.94 0.95
PoisonedRAG (White-Box) 0.91 0.98
6.3 LLM Agent
We also evaluate PoisonedRAG for LLM agents that interact with an external environment to obtain information for a task. We adopt the ReAct Agent framework [30], which combines reasoning and acting with LLMs. Given a question-answering task, the agent will perform a sequence of thought-action-observation steps to solve the task. The action space consists of two actions in our experiments: document retrieval and task finishing. For document retrieval, the agent will retrieve $k$ documents from a knowledge database (i.e., interacting with an environment to obtain information). For task finishing, the agent finishes the question-answering task and outputs the final answer. In each thought-action-observation step, the agent first generates a thought and an action. The thought provides a verbal reasoning processing for the next action (e.g., “I need to search Chicago Fire Season 4 and find how many episodes it has.”) to solve a task. Then, the agent takes the generated action (e.g., “Search [Chicago Fire Season 4]”) to obtain additional information (i.e., observation). Based on the additional information, the agent performs the next thought-action-observation step. This process is repeated until the task is finished (output final answer for the question-answering task) or a maximum number of steps is reached. We use the open-sourced code [30] in our experiment. We use the default setting of the previous evaluation and conduct the experiment on NQ, HotpotQA and MS-MARCO datasets. Our black-box attack achieves 0.72, 0.58, and 0.52 ASR, respectively.
@@ 7 Defenses
Many defenses [97, 98, 99, 100, 101, 102] were proposed to defend against data poisoning attacks that compromise the training dataset of a machine learning model. However, most of them are not applicable since PoisonedRAG does not compromise the training dataset of an LLM. Another defense is to (manually) check retrieved texts when observing generation error [103]. However, the generation error could also happen for many other reasons, making this solution time-consuming and less practical. Thus, we generalize some widely used defenses against attacks [44, 45, 46] to LLM to defend against PoisonedRAG.
@@ 7.1 Paraphrasing
Paraphrasing [44] was used to defend against prompt injection attacks [42, 48, 50, 51] and jailbreaking attacks [52, 53, 54, 55, 56, 57] to LLMs. We extend paraphrasing to defend against PoisonedRAG. In particular, given a text, the paraphrasing defense utilizes an LLM to paraphrase it. In our scenario, given a question, we use an LLM to paraphrase it before retrieving relevant texts from the knowledge database to generate an answer for it. Recall that PoisonedRAG crafts malicious texts such that they could be retrieved for a target question. For instance, in the black-box setting, PoisonedRAG prepends the target question to a text $I$ to craft a malicious text. In the white-box setting, PoisonedRAG optimizes a malicious text such that a retriever produces similar feature vectors for the malicious text and the target question. Our insight is that paraphrasing the target question would change its structure. For instance, when the target question is “Who is the CEO of OpenAI?”. The paraphrased question could be “Who holds the position of Chief Executive Officer at OpenAI?”. As a result, malicious texts may not be retrieved for the paraphrased target question. Note that we do not paraphrase texts in the knowledge database due to high computational costs.
@@
We conduct experiments to evaluate the effectiveness of paraphrasing defense. In particular, for each target question, we generate 5 paraphrased target questions using GPT-4, where the prompt can be found in Appendix H. For each paraphrased target question, we retrieve $k$ texts from the corrupted knowledge database (the malicious texts are crafted for the original target questions using PoisonedRAG). Then, we generate an answer for the paraphrased target question based on the $k$ retrieved texts. We adopt the same default setting as that in Section 5 (e.g., $k=5$ and 5 injected malicious texts for each target question). We report the ASR and F1-Score (note that Precision and Recall are the same as F1-Score under our default setting). ASR measures the fraction of paraphrased target questions whose answers are the corresponding attacker-chosen target answers. F1-Score is higher when more malicious texts designed for a target question are retrieved for the corresponding paraphrased target questions. Table 12 shows our experimental results. We find that PoisonedRAG could still achieve high ASRs and F1-Score, which means paraphrasing defense cannot effectively defend against PoisonedRAG.
@@ 7.2 Perplexity-based Detection
Perplexity (PPL) [104] is widely used to measure the quality of texts, which is also utilized to defend against attacks to LLMs [44, 45, 46]. A large perplexity of a text means it is of low quality. We utilize perplexity to detect malicious texts. For instance, in the white-box setting, PoisonedRAG utilizes adversarial attacks to craft malicious texts, which may influence the quality of malicious texts. Thus, a text with lower text quality (i.e., high perplexity) is more likely to be malicious. We calculate the perplexity for all clean texts in the database as well as all malicious texts crafted by PoisonedRAG. In our experiment, we use the cl100k_base model from OpenAI tiktoken [105] to calculate perplexity.
Figure 6 shows the ROC curve as well as AUC. We find that the false positive rate (FPR) is also very large when the true positive rate (TPR) is very large. This means a large fraction of clean texts are also detected as malicious texts when malicious texts are detected, i.e., the perplexity values of malicious texts are not statistically higher than those of clean texts, which means it is very challenging to detect malicious texts using perplexity. We suspect the reasons are as follows. Recall that each malicious text $P$ is the concatenation of $S$ and $I$, i.e., $P=S\oplus I$. The sub-text $I$ is generated by GPT-4, which is of high quality. For PoisonedRAG in the black-box setting, $S$ is the target question, which is a normal text. As a result, the text quality of the malicious text is normal. We find that the AUC of PoisonedRAG in the white-box setting is slightly larger than that in the black-box setting, which means the text quality is influenced by the optimization but not substantially.
Figure 6: The ROC curves for PPL detection defense. The dataset is NQ. The results for other two datasets are in Figures 7 and 8 in Appendix.
7.3 Duplicate Text Filtering
PoisonedRAG generates each malicious text independently in both black-box and white-box settings. As a result, it is possible that some malicious texts could be the same. Thus, we could filter those duplicate texts to defend against PoisonedRAG. We add experiments to filter duplicate texts under the default setting in Section 5. In particular, we calculate the hash value (using the SHA-256 hash function) for each text in a corrupted knowledge database and remove texts with the same hash value. Table 13 compares the ASR with and without defense. We find that the ASR is the same, which means duplicate text filtering cannot successfully filter malicious texts. The reason is that the sub-text $I$ (generated by GPT-4 in our experiment) in each malicious text is different, resulting in diverse malicious texts.
@@ 7.4 Knowledge Expansion
PoisonedRAG injects at most $N$ malicious texts into a knowledge database for each target question. Thus, if we retrieve $k$ texts, with $k>N$, then it is very likely that $k-N$ texts would be clean ones. This inspires us to retrieve more texts to defend against PoisonedRAG. We call this defense Knowledge Expansion. We conduct experiments under the default setting, where $N=5$. Figures 21, 22, 23 (in Appendix) shows the ASRs, Precision, Recall, and F1-Score for large $k$. We find that this defense still cannot completely defend against our PoisonedRAG even if $k=50$ (around 10% retrieved texts are malicious ones when injecting $N=5$ malicious texts for each target question). For instance, PoisonedRAG could still achieve 41% (black-box) and 43% (white-box) ASR on HotpotQA when $k=50$. Additionally, we find that ASR further increases as $N$ increases (shown in Figures 24, 25, 26 in Appendix), which means this defense is less effective when an attacker could inject more malicious texts into the knowledge database. We note that this defense also incurs large computation costs for an LLM to generate an answer due to the long context (caused by more retrieved texts).